# Quantization-free Lossy Image Compression

### v1

@@ Abstract
Lossy image compression is essential for efficient transmission and storage. Traditional compression methods mainly rely on discrete cosine transform (DCT) or singular value decomposition (SVD), both of which represent image data in continuous domains and therefore necessitate carefully designed quantizers. Notably, SVD-based methods are more sensitive to quantization errors than DCT-based methods like JPEG. To address this issue, we introduce a variant of integer matrix factorization (IMF) to develop a novel quantization-free lossy image compression method. IMF provides a low-rank representation of the image data as a product of two smaller factor matrices with bounded integer elements, thereby eliminating the need for quantization. We propose an efficient, provably convergent iterative algorithm for IMF using a block coordinate descent (BCD) scheme, with subproblems having closed-form solutions. Our experiments on the Kodak and CLIC 2024 datasets demonstrate that our IMF compression method consistently outperforms JPEG at low bit rates below 0.25 bits per pixel (bpp) and remains comparable at higher bit rates. We also assessed our method’s capability to preserve visual semantics by evaluating an ImageNet pre-trained classifier on compressed images. Remarkably, our method improved top-1 accuracy by over 5 percentage points compared to JPEG at bit rates under 0.25 bpp. The project is available at https://github.com/pashtari/lrf.
Keywords: Image Compression, Integer Matrix Factorization, Quantization-free Compression, Low-rank Approximation.
1.  Introduction
@@
Lossy image compression involves reducing the storage size of digital images by discarding some image data that are redundant or less perceptible to the human eye. This is crucial for efficiently storing and transmitting images, particularly in applications where bandwidth or storage resources are limited, such as web browsing, streaming, and mobile platforms. Lossy image compression methods enable adjusting the degree of compression, providing a selectable tradeoff between storage size and image quality. Widely used methods such as JPEG [wallace1991jpeg] and JPEG 2000 [skodras2001jpeg] follow the transform coding paradigm [goyal2001theoretical]. They use orthogonal linear transformations, such as discrete cosine transform (DCT) [ahmed1974discrete] and discrete wavelet transform (DWT) [antonini1992image], to decorrelate small image blocks. Since these transforms map image data into a continuous domain, quantization is necessary before coding into bytes. Unfortunately, as quantization errors can significantly degrade compression performance, the quantizers must be carefully crafted to minimize this impact, which further complicates codec design.
@@
Another promising paradigm relies on low-rank approximation techniques, with singular value decomposition (SVD) being a notable example. SVD is recognized as the deterministically optimal transform for energy compaction [andrews1976singular]. In practice, current SVD-based methods [andrews1976singular, prasantha2007image, hou2015sparse] can represent image data only with factors that contain floating-point elements, necessitating a quantization layer prior to any byte-level processing. However, this quantization often results in suboptimal compression performance, as SVD is more sensitive to quantization errors compared to transform-based methods like JPEG, especially at low bit rates.
@@
Motivated by this, we introduce a new variant of integer matrix factorization (IMF), and based on that, develop an effective quantization-free lossy image compression method. Our IMF formulation provides a low-rank representation of the image data as a product of two smaller factor matrices with bounded integer elements. Since we can directly store and losslessly process these integer matrices at the byte level, quantization is no longer needed, making IMF arguably better suited than SVD for image compression. Another advantage of IMF is that the reshaped factor matrices can be treated as 8-bit grayscale images, allowing any lossless image compression standard to be seamlessly integrated into the proposed framework. We propose an efficient iterative algorithm for IMF using a block coordinate descent (BCD) scheme, where each column of a factor matrix is taken as a block and updated one at a time using a closed-form solution.
@@
Our contributions are summarized as follows. We propose a new IMF formulation that enables quantization-free image compression. Moreover, we introduce an efficient algorithm for the IMF problem and prove its convergence. Finally, to the best of our knowledge, this work is the first effort to explore IMF for image compression, presenting the first algorithm based on a low-rank approach that significantly outperforms SVD and competes favorably with JPEG, particularly at low bit rates. Our method narrows the gap between factorization and quantization by integrating them into a single layer and optimizing the compression system.
2.  Related Work
@@
Transform coding is a widely used approach in lossy image compression, leveraging mathematical transforms to decorrelate pixel values and represent image data more compactly. One of the earliest and most influential methods is the discrete cosine transform (DCT) [ahmed1974discrete], used in JPEG [wallace1991jpeg], which converts image data into the frequency domain, prioritizing lower frequencies to retain perceptually significant information. The discrete wavelet transform (DWT) [antonini1992image], used in JPEG 2000 [skodras2001jpeg], offers improved performance by capturing both frequency and location information, leading to better handling of edges and textures [shapiro1993embedded]. More recently, the WebP [google2011webp] and HEIF [lainema2016hevc, hannuksela2015high] formats combine DCT and intra-frame prediction to achieve superior compression and quality compared to JPEG.
@@
[…]
@@
[…]
@@
[…]
3.  Method
3.1.  Overall Encoding Framework
The proposed compression method follows a transform coding paradigm, but it does not involve quantization. Figure 1 illustrates an overview of our encoding pipeline based on integer matrix factorization (IMF). The encoder accepts an RGB image with dimensions $H\times W$ and a color depth of 8 bits, represented by the tensor $\bm{\mathcal{X}}\in\{0,\ldots,255\}^{3\times H\times W}$. Each step of encoding is described in the following.
@@ 3.1.  Overall Encoding Framework
After chroma downsampling, we have three components: the luma component $\bm{Y}\in[0,255]^{H\times W}$ and the chroma components $\bm{C}_{B},\bm{C}_{R}\in[0,255]^{\frac{H}{2}\times\frac{W}{2}}$. Each of the matrices is split into non-overlapping $8\times 8$ patches. If a dimension of a matrix is not divisible by 8, the matrix is first padded to the nearest size divisible by 8 using reflection of the boundary values. These patches are then flattened into row vectors and stacked vertically to form matrices $\bm{X}_{Y}\in[0,255]^{\frac{HW}{64}\times 64}$, $\bm{X}_{C_{B}}\in[0,255]^{\frac{HW}{256}\times 64}$, and $\bm{X}_{C_{R}}\in[0,255]^{\frac{HW}{256}\times 64}$. Later, these matrices will be low-rank approximated using IMF. Note that this patchification technique differs from the block splitting in JPEG, where each block is subject to DCT individually and processed independently. This patchification technique not only captures the locality and spatial dependencies of neighboring pixels but also performs better when combined with the matrix decomposition approach for image compression.
@@
We now apply a low-rank approximation to the matrices $\bm{X}_{Y}$, $\bm{X}_{C_{B}}$, and $\bm{X}_{C_{R}}$, which is the core of our compression method that provides a lossy compressed representation of these matrices. The low-rank approximation [eckart1936approximation] aims to approximate a given matrix $\bm{X}\in\mathbb{R}^{M\times N}$ by
@@
where $\bm{U}\in\mathbb{R}^{M\times R}$ and $\bm{V}\in\mathbb{R}^{N\times R}$ are factor matrices (or simply factors), $R\leq\min(M,N)$ represents the rank, $U_{:r}$ and $V_{:r}$ represent the $r$-th columns of $\bm{U}$ and $\bm{V}$, respectively. We refer to $\bm{U}$ as the basis matrix and $\bm{V}$ as the coefficient matrix. By selecting a sufficiently small value for $R$, the factor matrices $\bm{U}$ and $\bm{V}$, with a combined total of $(M+N)R$ elements, offer a compact representation of the original matrix $\bm{X}$, which has $MN$ elements, capturing the most significant patterns in the image. Depending on the loss function used to measure the reconstruction error between $\bm{X}$ and the product $\bm{U}\bm{V}^{\mathsf{T}}$, as well as the constraints on the factor matrices $\bm{U}$ and $\bm{V}$, various formulations and variants have been proposed for different purposes [lee2000algorithms, ding2008convex, lin2005integer]. In Section 3.3, we introduce and elaborate on our variant, termed integer matrix factorization (IMF), and argue why it is well-suited and effective for image compression.
@@
IMF yields factor matrices $\bm{U}_{Y}\in\{0,\ldots,255\}^{\frac{HW}{64}\times R}$ and $\bm{V}_{Y}\in\{0,\ldots,255\}^{64\times R}$; $\bm{U}_{C_{B}}\in\{0,\ldots,255\}^{\frac{HW}{256}\times R}$ and $\bm{V}_{C_{B}}\in\{0,\ldots,255\}^{64\times R}$; and $\bm{U}_{C_{R}}\in\{0,\ldots,255\}^{\frac{HW}{256}\times R}$ and $\bm{V}_{C_{R}}\in\{0,\ldots,255\}^{64\times R}$ that correspond to $\bm{X}_{Y}$, $\bm{X}_{C_{B}}$, and $\bm{X}_{C_{R}}$. Since these matrices have integer elements, they can be directly encoded by any standard lossless data compression like zlib [deutsch1996zlib] without the need for a quantization step, which is commonly present in other lossy image compression methods and adds extra complications.
@@
As each channel of a factor map can be treated as an 8-bit grayscale image, we can encode it by any standard lossless image compression method such as PNG. For images with a resolution of $H,W\gg 64$, which are most common nowadays, the basis maps ($\bm{\mathcal{U}}$) are significantly larger than the coefficient maps ($\bm{\mathcal{V}}$), accounting for the majority of the storage space. Interestingly, in practice, the IMF basis maps turn out to be meaningful images, each capturing some visual semantic of the image (see Figure 2 for an example). Therefore, our IMF approach can effectively leverage the power of existing lossless image compression algorithms, offering a significant advantage over current methods. However, in this work, we take the first approach and use the zlib library [deutsch1996zlib] to encode factor matrices, creating a stand-alone codec that is independent from other image compression methods.
@@
Figure 2: The channels of IMF basis maps for the kodim23 image from Kodak. (a) shows the original image. The IMF basis maps corresponding to luma (b), blue-difference (c), and red-difference chroma (d) are shown. The channels of basis map with higher energy maintain the overall texture of the original image, where channels with lower energy focus more on subtle changes.
3.2.  Decoding
@@
3.3.  Integer Matrix Factorization (IMF)
The main building block of our method is integer matrix factorization (IMF), which is responsible for the lossy compression of matrices obtained through patchification. IMF can be framed as an optimization problem, aiming to minimize the reconstruction error between the original matrix $\bm{X}\in\mathbb{R}^{M\times N}$ and the product $\bm{U}\bm{V}^{\mathsf{T}}$, while ensuring that the elements of the factor matrices $\bm{U}$ and $\bm{V}$ are integers within a specified interval $[\alpha,\beta]$ with integer endpoints, i.e., $\alpha,\beta\in\mathbb{Z}$. Formally, the IMF problem can be expressed as:
@@ 3.3.  Integer Matrix Factorization (IMF)
where $\|\cdot\|_{\text{F}}$ denotes the Frobenius norm; $R\leq\min(M,N)$ represents the rank; and $\mathbb{Z}_{[\alpha,\beta]}\triangleq[\alpha,\beta]\cap\mathbb{Z}$ denotes the set of integers within $[\alpha,\beta]$. Without constraints on the factors, the problem would have an analytic solution through singular value decomposition (SVD), as addressed by the Eckart–Young–Mirsky theorem [eckart1936approximation]. If only a nonnegativity constraint were applied (without integrality), variations of nonnegative matrix factorization (NMF) would emerge [lee2000algorithms, gillis2020nonnegative]. The IMF problem (3.3) poses a challenging integer program, with finding its global minima known to be NP-hard [dong2018integer, van1981another]. Only a few iterative algorithms [dong2018integer, lin2005integer] have been proposed to find a “good solution” for some IMF variants in contexts other than image compression. In Section 3.4, we propose an efficient iterative algorithm for the IMF problem (3.3).
Algorithm 1 The proposed block coordinate descent (BCD) algorithm for IMF.
@@
The application of SVD and NMF in image compression is problematic mainly because the resulting factors contain continuous values that must be represented as arrays of floating-point numbers. This necessitates a quantization step that not only adds extra complications but also significantly degrades compression performance due to quantization errors (as demonstrated in Section 4). In contrast, our IMF formulation produces integer factor matrices that can be directly stored and losslessly processed without incurring roundoff errors. The reason for limiting the feasible region to $[\alpha,\beta]$ in our IMF formulation is to enable more compact storage of the factors using standard integral data types, such as int8 and int16, supported by programming languages. Given that the elements of the input matrix $\bm{X}$ are in $[0,255]$, we found the signed int8 type, which represents integers from -128 to 127, suitable for image compression applications. As a result, our IMF formulation is well-suited for image compression, effectively integrating the factorization and quantization steps into a single, efficient compression process.
3.4.  Block Coordinate Descent Scheme for IMF
We propose an efficient algorithm for IMF using the block coordinate descent (BCD) scheme (aka alternating optimization). The pseudocode is provided in Algorithm 1. Starting with some initial parameter values, this approach involves sequentially minimizing the cost function with respect to a single column of a factor at a time, while keeping the other columns of that factor and the entire other factor fixed. This process is repeated until a stopping criterion is met, such as when the change in the cost function value falls below a predefined threshold or the maximum number of iterations is reached. Formally, this involves solving one of the following subproblems at a time:

### v2

# Quantization-aware Matrix Factorization for Low Bit Rate Image Compression

@@ Abstract
Lossy image compression is essential for efficient transmission and storage. Traditional compression methods mainly rely on discrete cosine transform (DCT) or singular value decomposition (SVD), both of which represent image data in continuous domains and, therefore, necessitate carefully designed quantizers. Notably, these methods consider quantization as a separate step, where quantization errors cannot be incorporated into the compression process. The sensitivity of these methods, especially SVD-based ones, to quantization errors significantly degrades reconstruction quality. To address this issue, we introduce a quantization-aware matrix factorization (QMF) to develop a novel lossy image compression method. QMF provides a low-rank representation of the image data as a product of two smaller factor matrices, with elements constrained to bounded integer values, thereby effectively integrating quantization with low-rank approximation. We propose an efficient, provably convergent iterative algorithm for QMF using a block coordinate descent (BCD) scheme, with subproblems having closed-form solutions. Our experiments on the Kodak and CLIC 2024 datasets demonstrate that our QMF compression method consistently outperforms JPEG at low bit rates below 0.25 bits per pixel (bpp) and remains comparable at higher bit rates. We also assessed our method’s capability to preserve visual semantics by evaluating an ImageNet pre-trained classifier on compressed images. Remarkably, our method improved top-1 accuracy by over 5 percentage points compared to JPEG at bit rates under 0.25 bpp. The project is available at https://github.com/pashtari/lrf.
keywords:
Matrix Factorization, Low-rank Approximation, Quantization, Image Compression.
journal: Information Sciences
[…]
[KUL]organization=Department of Electrical Engineering (ESAT), STADIUS Center, KU Leuven, country=Belgium
[…]
[Std]organization=Stanford Center for Biomedical Informatics Research, Stanford University, city=CA, country=USA
1 Introduction
@@
Lossy image compression involves reducing the storage size of digital images by discarding some image data that are redundant or less perceptible to the human eye. This is crucial for efficiently storing and transmitting images, particularly in applications where bandwidth or storage resources are limited, such as web browsing, streaming, and mobile platforms. Lossy image compression methods enable adjusting the degree of compression, providing a selectable tradeoff between storage size and image quality. Widely used methods such as JPEG [1] and JPEG 2000 [2] follow the transform coding paradigm [3]. They use orthogonal linear transformations, such as discrete cosine transform (DCT) [4] and discrete wavelet transform (DWT) [5], to decorrelate small image blocks. Since these transforms map image data into a continuous domain, quantization is necessary before coding into bytes. Unfortunately, as quantization errors can significantly degrade compression performance, the quantizers must be carefully crafted to minimize this impact, which further complicates codec design.
@@
Another promising paradigm relies on low-rank approximation techniques, with singular value decomposition (SVD) being a notable example. SVD is recognized as the deterministically optimal transform for energy compaction [6]. In practice, current SVD-based methods [6, 7, 8] can represent image data only with factors that contain floating-point elements, necessitating a quantization step prior to any byte-level processing. The quantization step often introduces errors, which result in suboptimal compression performance.
@@
Motivated by this, we introduce quantization-aware matrix factorization (QMF) and, based on it, develop an effective lossy image compression method. Unlike traditional compression methods, the proposed approach integrates quantization into the optimization process rather than treating it as a separate step before byte-level processing. Our QMF formulation provides a low-rank representation of image data as the product of two smaller factor matrices. The quantization is integrated via introducing constraints in the optimization process, where the elements of the factor matrices are constrained to bounded integer values. These elements, with discrete values represented as bounded integers, can be directly stored using standard integral data types—such as int8 and int16 supported by programming languages—and losslessly processed, making QMF arguably better suited than SVD for image compression. Another advantage of QMF is that the reshaped factor matrices can be treated as 8-bit grayscale images, allowing any lossless image compression standard to be seamlessly integrated into the proposed framework. We propose an efficient iterative algorithm for QMF using a block coordinate descent (BCD) scheme, where each column of a factor matrix is taken as a block and updated one at a time using a closed-form solution.
@@
Our contributions are summarized as follows. We propose a novel optimization framework that enables the integration of quantization and low-rank approximation for image compression. Moreover, we introduce an efficient algorithm for the QMF problem and prove its convergence. Finally, to the best of our knowledge, this work is the first effort to explore QMF for image compression, presenting the first algorithm based on a low-rank approach that significantly outperforms SVD and competes favorably with JPEG, particularly at low bit rates. Our method narrows the gap between factorization and quantization by integrating them into a single layer and optimizing the compression system.
2 Related Work
@@
Transform coding is a widely used approach in lossy image compression, leveraging mathematical transforms to decorrelate pixel values and represent image data more compactly. One of the earliest and most influential methods is the discrete cosine transform (DCT) [4], used in JPEG [1], which converts image data into the frequency domain, prioritizing lower frequencies to retain perceptually significant information. The discrete wavelet transform (DWT) [5], used in JPEG 2000 [2], offers improved performance by capturing both frequency and location information, leading to better handling of edges and textures [9]. More recently, the WebP [10] and HEIF [11, 12] formats combine DCT and intra-frame prediction to achieve superior compression and quality compared to JPEG.
@@
Recently, learned image compression (LIC) has gained attention for potentially outperforming traditional methods by leveraging deep neural networks. Ballé et al. [13] pioneered this area with an end-to-end trainable convolutional neural network based on variational autoencoders. Cheng et al. [14] incorporated a simplified attention module and discretized Gaussian mixture likelihoods for achieving a more accurate and flexible entropy model. Liu et al. [15] combined transformers and CNNs to exploit the local modeling ability of convolutions and the global modeling ability of the attention mechanism. Yang and Mandt [16] introduced diffusion models into LIC, using a denoising decoder to iteratively reconstruct a compressed image. Despite these advancements, the high computational complexity of LIC methods remains a significant limitation, particularly for real-time applications and resource-constrained environments.
@@
Low-rank approximation can provide a compact representation by decomposing image data into smaller components. Notably, truncated singular value decomposition (tSVD) is a classical technique that decomposes images into singular values and vectors, retaining only the most significant components to achieve compression [6, 7]. Hou et al. [8] proposed sparse low-rank matrix approximation (SLRMA) for data compression, which is able to explore both the intra- and inter-coherence of data samples simultaneously from the perspective of optimization and transformation. More recently, Yuan and Haimi-Cohen [17] introduced a graph-based low-rank regularization to reduce compression artifacts near block boundaries at low bit rates.
@@
There are applications where meaningful representation of data as discrete factor matrices is crucial. While typical low-rank techniques like SVD and nonnegative matrix factorization (NMF) are inappropriate for such applications, integer matrix factorization (IMF) ensures the integrality of factors to achieve this goal. Lin et al. [18] investigates IMF to effectively handle discrete data matrices for cluster analysis and pattern discovery. Dong et al. [19] introduce an alternative least squares method for IMF, verifying its effectiveness with some data mining applications. However, the application of IMF in image compression remains unexplored.
While existing IMF methods generally constrain factor elements to the entire set of integer values, we propose quantized matrix factorization (QMF), which minimizes the objective function over a bounded interval of integers, thereby modeling a uniformly quantized domain. Furthermore, we introduce a block coordinate descent (BCD)-based algorithm to solve the QMF problem, which is both computationally efficient and provably convergent. This work investigates the potential of QMF for image compression, arguing that it can serve as a powerful tool for this purpose.
3 Method
3.1 Overall Encoding Framework
The proposed compression method follows a transform coding paradigm, but it does not involve a separate quantization step. Figure 1 illustrates an overview of our encoding pipeline based on quantization-aware matrix factorization (QMF). The encoder accepts an RGB image with dimensions $H\times W$ and a color depth of 8 bits, represented by the tensor $\bm{\mathcal{X}}\in\{0,\ldots,255\}^{3\times H\times W}$. Each step of encoding is described in the following.
@@ 3.1 Overall Encoding Framework
After chroma downsampling, we have three components: the luma component $\bm{Y}\in[0,255]^{H\times W}$ and the chroma components $\bm{C}_{B},\bm{C}_{R}\in[0,255]^{\frac{H}{2}\times\frac{W}{2}}$. Each of the matrices is split into non-overlapping $8\times 8$ patches. If a dimension of a matrix is not divisible by 8, the matrix is first padded to the nearest size divisible by 8 using reflection of the boundary values. These patches are then flattened into row vectors and stacked vertically to form matrices $\bm{X}_{Y}\in[0,255]^{\frac{HW}{64}\times 64}$, $\bm{X}_{C_{B}}\in[0,255]^{\frac{HW}{256}\times 64}$, and $\bm{X}_{C_{R}}\in[0,255]^{\frac{HW}{256}\times 64}$. Later, these matrices will be low-rank approximated using QMF. Note that this patchification technique differs from the block splitting in JPEG, where each block is subject to DCT individually and processed independently. This patchification technique not only captures the locality and spatial dependencies of neighboring pixels but also performs better when combined with the matrix decomposition approach for image compression.
@@
We now apply a low-rank approximation to the matrices $\bm{X}_{Y}$, $\bm{X}_{C_{B}}$, and $\bm{X}_{C_{R}}$, which is the core of our compression method that provides a lossy compressed representation of these matrices. The low-rank approximation [20] aims to approximate a given matrix $\bm{X}\in\mathbb{R}^{M\times N}$ by
@@
where $\bm{U}\in\mathbb{R}^{M\times R}$ and $\bm{V}\in\mathbb{R}^{N\times R}$ are factor matrices (or simply factors), $R\leq\min(M,N)$ represents the rank, $U_{:r}$ and $V_{:r}$ represent the $r$-th columns of $\bm{U}$ and $\bm{V}$, respectively. We refer to $\bm{U}$ as the basis matrix and $\bm{V}$ as the coefficient matrix. By selecting a sufficiently small value for $R$, the factor matrices $\bm{U}$ and $\bm{V}$, with a combined total of $(M+N)R$ elements, offer a compact representation of the original matrix $\bm{X}$, which has $MN$ elements, capturing the most significant patterns in the image. Depending on the loss function used to measure the reconstruction error between $\bm{X}$ and the product $\bm{U}\bm{V}^{\mathsf{T}}$, as well as the constraints on the factor matrices $\bm{U}$ and $\bm{V}$, various formulations and variants have been proposed for different purposes [21, 22, 18]. In Section 3.3, we introduce and elaborate on our variant, termed quantization-aware matrix factorization (QMF), and argue why it is well-suited and effective for image compression.
@@
QMF yields factor matrices $\bm{U}_{Y}\in\{0,\ldots,255\}^{\frac{HW}{64}\times R}$ and $\bm{V}_{Y}\in\{~{}0,~{}~{}~{}~{}~{}\ldots\\ ,255\}^{64\times R}$; $\bm{U}_{C_{B}}\in\{0,\ldots,255\}^{\frac{HW}{256}\times R}$ and $\bm{V}_{C_{B}}\in\{0,\ldots,255\}^{64\times R}$; and $\bm{U}_{C_{R}}\in\{0,\ldots,255\}^{\frac{HW}{256}\times R}$ and $\bm{V}_{C_{R}}\in\{0,\ldots,255\}^{64\times R}$ that correspond to $\bm{X}_{Y}$, $\bm{X}_{C_{B}}$, and $\bm{X}_{C_{R}}$. Since these matrices have elements constrained to integer values (allowing seamless integration of quantization with their optimization process), they can be directly encoded using any standard lossless data compression method, such as zlib [23]. In contrast, other lossy image compression methods typically require a separate quantization step, introducing errors that cannot be incorporated or considered during the compression process.
@@
As each channel of a factor map can be treated as an 8-bit grayscale image, we can encode it by any standard lossless image compression method such as PNG. For images with a resolution of $H,W\gg 64$, which are most common nowadays, the basis maps ($\bm{\mathcal{U}}$) are significantly larger than the coefficient maps ($\bm{\mathcal{V}}$), accounting for the majority of the storage space. Interestingly, in practice, the QMF basis maps turn out to be meaningful images, each capturing some visual semantic of the image (see Figure 2 for an example). Therefore, our QMF approach can effectively leverage the power of existing lossless image compression algorithms, offering a significant advantage over current methods. However, in this work, we take the first approach and use the zlib library [23] to encode factor matrices, creating a stand-alone codec that is independent from other image compression methods.
@@
Figure 2: The channels of QMF basis maps for the kodim23 image from Kodak. (a) shows the original image. The QMF basis maps corresponding to luma (b), blue-difference (c), and red-difference chroma (d) are shown. The channels of basis map with higher energy maintain the overall texture of the original image, where channels with lower energy focus more on subtle changes.
3.2 Decoding
@@
3.3 Quantization-aware Matrix Factorization (QMF)
The main building block of our method is quantization-aware matrix factorization (QMF), which is responsible for the lossy compression of matrices obtained through patchification. QMF can be framed as an optimization problem, aiming to minimize the reconstruction error between the original matrix $\bm{X}\in\mathbb{R}^{M\times N}$ and the product $\bm{U}\bm{V}^{\mathsf{T}}$, while ensuring, as an integrated quantization step, that the elements of the factor matrices $\bm{U}$ and $\bm{V}$ are integers within a specified interval $[\alpha,\beta]$ with integer endpoints, i.e., $\alpha,\beta\in\mathbb{Z}$. Formally, the QMF problem can be expressed as:
@@ 3.3 Quantization-aware Matrix Factorization (QMF)
where $\|\cdot\|_{\text{F}}$ denotes the Frobenius norm; $R\leq\min(M,N)$ represents the rank; and $\mathbb{Z}_{[\alpha,\beta]}\triangleq[\alpha,\beta]\cap\mathbb{Z}$ denotes the set of integers within $[\alpha,\beta]$. Without constraints on the factors, the problem would have an analytic solution through singular value decomposition (SVD), as addressed by the Eckart–Young–Mirsky theorem [20]. If only a nonnegativity constraint were applied (without integrality), variations of nonnegative matrix factorization (NMF) would emerge [21, 24]. The QMF problem (3.3) poses a challenging integer program, with finding its global minima known to be NP-hard [19, 25]. Only a few iterative algorithms [19, 18] have been proposed to find a “good solution” for some QMF variants in contexts other than image compression. In Section 3.4, we propose an efficient iterative algorithm for the QMF problem (3.3).
Algorithm 1 The proposed block coordinate descent (BCD) algorithm for QMF.
@@
The existing lossy image compression methods based on SVD and NMF approach the problem as an optimization task, followed by a separate quantization step. The optimization focuses on finding factors that minimize the reconstruction error. Before any byte-level processing, a quantization step is applied to project the floating-point elements of the resulting factors onto a set of discrete values. However, because the quantization is performed separately from the optimization, the quantization errors cannot be incorporated into the compression process. This separation leads to suboptimal compression performance (as demonstrated in Section 4) and additional complications in designing quantizers. In contrast, our QMF formulation, through the constrained optimization in (3.3), produces integer factor matrices that minimize the reconstruction error while ensuring that all elements are discrete. These integer factor matrices can be directly stored and processed losslessly without introducing roundoff errors. The reason for limiting the feasible region to $[\alpha,\beta]$ in our QMF formulation is to enable more compact storage of the factors using standard integral data types, such as int8 and int16, supported by programming languages. Given that the elements of the input matrix $\bm{X}$ are in $[0,255]$, we found the signed int8 type, which represents integers from -128 to 127, suitable for image compression applications. As a result, our QMF formulation is well-suited for image compression, effectively integrating the factorization and quantization steps into a single, efficient compression process.
3.4 Block Coordinate Descent Scheme for QMF
We propose an efficient algorithm for QMF using the block coordinate descent (BCD) scheme (aka alternating optimization). The pseudocode is provided in Algorithm 1. Starting with some initial parameter values, this approach involves sequentially minimizing the cost function with respect to a single column of a factor at a time, while keeping the other columns of that factor and the entire other factor fixed. This process is repeated until a stopping criterion is met, such as when the change in the cost function value falls below a predefined threshold or the maximum number of iterations is reached. Formally, this involves solving one of the following subproblems at a time: